A stone is thrown upwards with an accelaration of 10 m/sec.sq.If its velocity is 10m/sec, then find the distance and taken to reach the ground.


Please suggest the right method to do it


thanks

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Neel

stone will go upwards  and at its maximum height its velocity will be zero

apply v2= u2 + 2as

put,v=0, u= 10, a=g=-10 m/s2

s=?

0=100-20*s

it gives s= 5m

it will cover total distance of 2*5=10m (up and down)

for calculation of time use s=ut + 0.5at2

remember s in these equtions is displacment , when stone go up and then come down total displacement will be zero.

s=0, u= 10, a=g=-10

put these values it gives t=2 sec.

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.


AKASH GOYAL
AskiitiansExpert-IITD